OpenStudy (anonymous):

Please help asap A biologist is comparing the growth of a population of flies per week to the number of flies a bullfrog will consume per week. She has devised an equation to solve for which day (x) the bullfrog would be able to eat the entire population. The equation is 2x = 3x - 1. However, she has observed that the bullfrog cannot eat more than seven flies in one week. Explain to the biologist how she can solve this on a graph using a system of equations. Identify any possible constraints to the situation.

OpenStudy (campbell_st):

plot the lines y = 2x and y = 3x - 1 the point where they intersect is the day where the bullfrogs eat all the flies... and it will also give the number of flies eaten

OpenStudy (anonymous):

its (1,2) but what about the constraints. I dont get that @campbell_st

OpenStudy (campbell_st):

so plot the lines... I don't know which line represents each animal... well for me the constraint is that it's a 1st quadrant graph, so x >= 0 and the line y = 3x - 1has a y-intercept of y = -1 so it would look like 

OpenStudy (anonymous):

its 2^x

OpenStudy (campbell_st):

the reason x > = 0 is that time is only positive and perhaps y > = 0 as you can't have a negative population

OpenStudy (campbell_st):

ok... so are the equations both exponential...?

OpenStudy (anonymous):

oh okay and they cant eat neagtive flies so x and y have to be 0 or greater than 0 right

OpenStudy (campbell_st):

\[2^x = 3^{x - 1}\]

OpenStudy (anonymous):

no its 2^x and 3x-1

OpenStudy (anonymous):

only 2^x is exponential

OpenStudy (campbell_st):

ok... so same thing applies... 

OpenStudy (anonymous):

they intersect at (1,2) what would that stand for though?

OpenStudy (anonymous):

geogebra says its 1,2

OpenStudy (campbell_st):

oops they do... sorry

OpenStudy (anonymous):

okay but what would that stand for

OpenStudy (campbell_st):

so the point in intersection is the day when the populations are equal in number.

OpenStudy (anonymous):

x=domain what does y=?

OpenStudy (campbell_st):

well there are 2 points of interection.. (1, 2) and (3, 8)

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

so are those the two days they are equal or eat the most?

OpenStudy (campbell_st):

well reading the question those are the days the bullfrogs eat the entire population... but on day 3 the population is 8... and they can only eat 7... which is the max of the range... so the domain is restricted to 7 = 3x - 1 or x = 8/3 which is \[0 \le x \le \frac{8}{3}\]

OpenStudy (campbell_st):

that is assuming the y = 3x -1 represents the bullfrogs...

OpenStudy (anonymous):

But you know how they say x is the domain, what would y be

OpenStudy (anonymous):

and wouldnt that be y since x is the day?

OpenStudy (campbell_st):

to me, y is the number of flies in the population...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

can u help me with one more?